Find the location of the radial node for an electron in a 3p orbital of a hydrogen atom (in meters)
Answer #1

There are two types of nodes:A radial node is a sphere that occurs when the radial wavefunction for an atomic orbital is equal to zero or changes sign.

angular node is a flat plane.

Radial nodes occur as the principle quantum number (n) increases and the number of radial nodes depends on the principle quantum number (n) and the number of angular nodes (l). The total number of nodes is calculated as:

  Total Nodes=n-1

and

    Radial Nodes=n-l-1

where n is the principle quantum number,l is the azimuthal quantum number

for 3rd shell,total number of nodes=n-1.i.e 3-1=2

for 3s orbital l=0 so there are 0 angular nodes and 2 radial nodes

   Radial Nodes=n-l-1

here,l=0 thus 3-1=2

for 3p orbital l=1 so no.of radial nodes-n-l-1.i.e 3-1-1=1

therefore there is one radial node and one angular node

for 3d orbital l=2 so there are 2 angular nodes and 0 radial nodes.